OpenStudy (anonymous):

What is the Value of A in the equation? 7x9=(7x10)-(7xA)

OpenStudy (isaiah.feynman):

We dunno x and we dunno A, how do we solve one equation with two unknowns?

OpenStudy (anonymous):

it's multiflicationI guess I could've put 7*9=(7*10)-(7*A)

OpenStudy (anonymous):

multiplication

OpenStudy (isaiah.feynman):

63=70-7A, So do you know what next to do?

OpenStudy (anonymous):

I'm pelletty at Distributive Property!...

OpenStudy (anonymous):

crappy*

OpenStudy (anonymous):

0?

OpenStudy (isaiah.feynman):

No need to distribute anything here. How do you get 7A isolated?

OpenStudy (anonymous):

no clue....o.o

OpenStudy (isaiah.feynman):

Subtract 70 from both sides to get -7=-7A. Can you solve for A from here?

OpenStudy (anonymous):

-7?

OpenStudy (isaiah.feynman):

Lol are you guessing?

OpenStudy (anonymous):

0? and yes cause I have no idea what you're talking about.....

OpenStudy (isaiah.feynman):

Well that's heartbreaking, I dunno what happens in your class, but it seems I should teach how to solve this little equations so they will be at your mercy.

OpenStudy (anonymous):

Okay? This is the only math question that I have that is absolutely pissing me off!....

OpenStudy (isaiah.feynman):

Well how about a little teaching, so you don't get this question again in the future to piss you off.

OpenStudy (anonymous):

?

OpenStudy (isaiah.feynman):

I want to show you have to solve this little equations.

OpenStudy (anonymous):

I am a slow and terrible learner but okay...

OpenStudy (isaiah.feynman):

What is x/x?

OpenStudy (anonymous):

x/x????????

OpenStudy (isaiah.feynman):

Yes, thats basic, you must know how to solve that in order to solve complex ones like...2=x/2

OpenStudy (anonymous):

I can solve this yet I can't solve the question I've just asked.... something is wrong here..... \[\frac{ x }{ 8 }+12=16\]

OpenStudy (isaiah.feynman):

Open study can't show the equatio well.

OpenStudy (anonymous):

x over 8 plus 12 equals 16

OpenStudy (isaiah.feynman):



OpenStudy (anonymous):

x=36

OpenStudy (isaiah.feynman):

32 actually.

OpenStudy (anonymous):

ops I was looking at wrong equation I had on my paper lol :P

OpenStudy (anonymous):

haha my bad.. I was looking at 48-6*2=36

OpenStudy (isaiah.feynman):

Okay, it seems you can solve this equations. So back to the one at hand.